The Michaelis-Menten equation is often used to describe the kinetic characteristics of an enzyme-catalyzed reaction. S Where v is the velocity or rate, Vmax is the maximum velocity, Km is the +IST Michaelis- Menten constant, and I5 s the substrate concentration. K + S v (uM/min) a) A graph of the Michaelis-Menten equation is a plot of a reaction's initial velocity (Vo) at different substrate concentrations ([S]) 300 Vmax 250 1/2 Vmax First, move the line labeled "Vmax to a position that represents the maximum velocity of the enzyme 225 Next, move the line labeled "1/2 Vmax" to its correct position. 150 Km Michaeli ten Curve Then, move the line labeled "Km" to its correct position. 100 01° 20 30 40 50 60 70 80 90 100 [SI (AM) below... b) Estimate the values for these important constants: ⓧGive Up & View Solution-e,Check Answer O Previous Next Exit
correct position. 10) 20 30 40-50,,.60 70 80 90 100 [S) (uM) Continues below b) Estimate the values for these important constants: Number Number HM/ min
Answer #1

According to the michaelis menten graph, the Vmax is the highest value of the curve, which is 175 uM / min, the Vmax line is located at 175 uM / min.

The V 1/2 is then calculated:

V 1/2 = Vmax / 2 = 175/2 = 87.5 uM / min

The V 1/2 line is located at 87.5 uM / min

Km is read on the horizontal axis, precisely corresponding to the value of V 1/2 and Km is: 11 uM.

If you liked the answer please rate it in a positive way, you would help me a lot, thank you.